0 14 + x2 2. (3 points) Solve the initial value problem xy–3y = x?, y(i)=
Answer #1

Given xey-3y = x² gely!= 34) = * ? y- 34 XХ ts a y + ( 22 ) first order linear diferential the form. equation, Of dy de tpy- a yt py=Here P =-3 Q = x X Integrating factor (If) = el de S-3 da =é -3 Is (x) e es 37 3 If X Gueneral Solution y (If) = fa (If) dx +- 2 1 1 2 de tc - 2+1 x +C - 24 y 3 x 1 / 2 tc y = -x3 + Cx² 4(x)= -&2&ca 3 Gwen initial Condition у (i) - y (1) = - (1) 2 +